OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#4652 — var-1-6k/PLIX
Incident Report for Network & Infrastructure
Resolved
We have got a problem on the port-channel Amsterdam<>Warsaw and that towards PLIX. Probably, these problems are caused by a hardware problem on the card #3 of our network at Warsaw. We will restart this card.

Update(s):

Date: 2010-10-06 22:14:07 UTC
We re-pass the PLIX to a normal priority.



Date: 2010-10-06 16:27:16 UTC
The fiber was cleaned and the fiber connections were verified. We will reinstate the 2nd port.



Date: 2010-10-03 21:57:34 UTC
The PLIX team has changed the Xenpack on our server.
The problem is not fixed but there is less errors

We are going to remove the second 10G from production
in order to fix the problem.

Date: 2010-10-03 17:33:56 UTC
It seems that there are errors in the PLIX on the reception of packets
coming from OVH . We actually don't see any errors.

Well , we're on the good track ...

Port Rx Error Monitor Sun Oct 3
16:32:49 2010
Port Link Rx Rx Rx Rx Rx Rx
Rx
State Crc Over Under Frag Jabber Align
Lost
================================================================================
PLIX-OVH> A 482794159 4953587 0 0 0 0
0

Port Tx Error Monitor Sun Oct 3
16:33:26 2010
Port Link Tx Tx Tx Tx Tx
Tx
State Coll Late coll Deferred Errors Lost
Parity
================================================================================
PLIX-OVH> A 0 0 0 0 0
0

--------------------------------------


Date: 2010-10-03 15:14:52 UTC
We remark the same problem on the card #3 even after
the card change.

We have cut off the card #3.

We will update the router.

Date: 2010-10-03 15:13:17 UTC
The replacement was done. The service is again optimal.

Date: 2010-09-30 10:09:18 UTC
We will proceed to the replacement of the card in about 15min.

Date: 2010-09-29 15:10:05 UTC
We have deactivated the card #3 awaiting to receive
the new one. We have diminished the BGP weight of
PLIX and we rerouted the AS via DECIX or AMSIX.

Date: 2010-09-29 10:39:15 UTC
The problem continues. We will reload the card. If this continues,
we will change it.

Date: 2010-09-29 10:38:29 UTC
The traffic is well balanced again. We will keep an eye on this card.
If the problem starts again, we will change it.
Posted Sep 29, 2010 - 10:36 UTC